# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v1

@@ Abstract
This Letter systematizes the study of ALT phases in strongly interacting systems. Specifically, we focus on localized topological phases satisfying a condition we refer to as short-ranged entanglement. We will argue that instead of eigenstate order, the right way to characterize all topological localized phases with short-ranged entanglement is via quantum cellular automata (QCA) [58, 59, 60, 56]—the locality preserving unitaries which map physical degrees of freedom to local integrals of motion (LIOMs) (Fig. 1(a)).
As QCA in spin systems have been classified in one [58] and two [61] dimensions, this correspondence allows for a classification of ALT phases in low dimensional systems. The same formalism extends to periodically (Floquet [62]) and quasiperiodically driven ALT phases [31], which are classified by loops or tori of QCA, respectively. Thus, we can treat static and $n$-tone-driven ALT phases in $d$ dimensions on the same footing. We also considerably develop the study of the topology of QCA, which allows us to classify all ALT phases with $d-n\leq 2$ (Fig. 1(b)).
@@
We say two MBL Hamiltonians $H_{0}$ and $H_{1}$ are in the same localized topological phase if there is a continuous path of MBL Hamiltonians $H_{t}$ ($t\in[0,1]$) which connects them. Ignoring long-range resonances in the spectrum of $H_{t}$, it is possible to construct a continuous path of locality preserving unitaries $V_{t}$ (that is, $V_{t}AV_{t}^{\dagger}$ is localized near the support of $A$ for all operators $A$) such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$ [84, 85, 86, 87, 88, 89]. (Indeed, this unitary simply follows the adiabatic evolution of the many-body eigenstates.) Including long-range resonances prevents an adiabatic construction of $V_{t}$ [90], but it is expected that sufficiently dilute resonances can be incorporated such that there still exists a $V_{1}$ that conjugates LIOMs of $H_{0}$ into those of $H_{1}$ and is continuously connected to the identity [9]. Thus, if $H_{0}$ and $H_{1}$ are in the same localized phase, their LIOMs are related by a continuous path of locality-preserving unitary operators. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@
$\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}\implies\pi_{n-1}(\mathrm{QCA}_{d-1})\cong\pi_{n}(\mathrm{QCA}_{d})$ (5)
for all $n\geq 1$. The base point for the loops in $\Omega\mathrm{QCA}_{d}$ is the identity.
@@
Equation (5) is interpreted through a picture of pumping [95, 104, 103]. Taking a loop $\gamma\in\Omega\mathrm{QCA}_{d}$ and restricting it to open boundary conditions, the action on the bulk must be trivial at the end of the loop, due to the finite range of the QCA. However, on the $(d-1)$-dimensional edge, there may be a nontrivial action by some $v\in\mathrm{QCA}_{d-1}$ [23]. We say that $\gamma$ pumps $v$ across the system, and Eq. (5) hypothesizes that the homotopy structure of the loop space $\Omega\mathrm{QCA}_{d}$ is completely characterized by the pumped QCA $v$.
@@
The conjectured classification of symmetry-enriched invertible states can be stated in terms of just the space $\mathrm{inv}_{d}$ and the abstract symmetry group $G$ [96, 66, 95, 102, 113, 103, 104]. However, the classification of symmetric QCA depends not just on $G$, but on the specific representation $\rho$ of the symmetry action [93, 114]. We compute the classification of symmetric QCA, and thus SALT phases, with $d-n\leq 1$ and any on-site unitary representation of any group in Appendix D. The result closely mirrors Tab. 1, except that rational numbers $p/q\in{\mathbb{Q}}$ are replaced with formal ratios of representations, $\rho/\sigma$. Addition and multiplication for $\rho/\sigma$ is defined through the same rules as for ratios of integers, but replacing “$+$” with $\oplus$” and “$\times$” with $\otimes$”. In an appropriate coincidence, this mathematical structure is known as an algebraic localization.
To define the algebraic localization more precisely, we must specify the set of ancillae representations,
@@
which is closed under tensor product—if $\rho_{1}$ and $\rho_{2}$ can be added then $\rho_{1}\otimes\rho_{2}$ corresponds to adding them both. The formal fractions $\rho/\sigma$ of the algebraic localization with denominators in the set $\mathcal{R}$ belong to the quotient
@@
where $R_{G}$ is the representation ring of $G$, and $\simeq$” is unitary equivalence of representations.
@@
Models.—While the swindle map (Eq. (8), Fig. 2) was introduced as an abstract tool, it also defines a class of soluble models of $n$-tone-driven (S)ALT phases given any $(n-1)$-parameter family of QCA in one dimension lower. As these models are finite depth circuits, they are also well suited for realization in quantum simulator experiments [63, 64, 65] (Fig. 3).
@@
The general relationship between (S)ALT phases and QCA we have demonstrated, as well as the $\Omega$-spectrum property which relates static (S)ALT phases to driven phases, extends to fermionic systems essentially without modification (Appendix A-B[56]. The $\Omega$-spectrum structure of the classification can also be exploited to classify SALT phases with lattice symmetries [116, 103, 104]. The more challenging direction of generalization is to include topological order (long-ranged entanglement), where the correspondence with QCA breaks down [20]. This case is of particular interest, given recent proposals of using Floquet driving of topologically ordered localized models to perform quantum computation [117, 118, 119, 120]. The notion of measurement QCA introduced by Ref. [121] may provide a route towards classifying localized phases with topological order, which we will explore in future work.
The authors are particularly grateful to R. Kobayashi for collaboration during the early stages of this project. The authors also thank M. Barkeshli, A. Chandran, L. Fidkowski, S. Pace, R. Thorngren, Y.-T. Tu, B. Ware, and C. Zhang for useful discussions. This work is supported by the Laboratory for Physical Sciences (D.M.L.). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities (D.V.E.).
@@ A.1 Setup
(where $B_{r}(j)$ is the ball of radius $r$ around $j$) such that $[\tau^{z}_{j},H]=[\tau^{z}_{j},\tau^{z}_{j^{\prime}}]=0$. Further, the set of LIOMs must be maximal, such that a local operator $A$ commutes with all LIOMs only if $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$.
@@
This definition should be compared to [56, Definition II.3]. One can also compare this notion of “short-ranged entanglement” for Hamiltonians to the usual notion of short-ranged entanglement for states. A state $|\psi\rangle$ may be said to be short-range entangled if, for any other state $|\psi^{\prime}\rangle$ which differs from $|\psi\rangle$ only in some ball $B_{r}(x)$ ($\langle\psi|A|\psi\rangle=\langle\psi^{\prime}|A|\psi^{\prime}\rangle$ for any local $A$ not supported in $B_{r}(x)$) there is a local unitary $\tau^{x}_{\psi\psi^{\prime}}$ supported in $B_{r}(x)$ such that $\tau^{x}_{\psi\psi^{\prime}}|\psi^{\prime}\rangle=|\psi\rangle$. Any eigenstate of an $H$ satisfying Definition A.2 is short-range entangled in this sense (the required $\tau^{x}_{\psi\psi^{\prime}}$ can be constructed from the $\tau^{x}_{j}$ operators), but the converse—that $\tau^{x}_{j}$ operators exist for any MBL Hamiltonian $H$ whose eigenstates are all short-range entangled—is not obvious.
@@ A.2 Existence
is isomorphic to a matrix algebra. This implies that there is an isomorphism between $\mathcal{T}_{j}$ and $\mathcal{O}_{k}$ whenever they have the same dimension. (All isomorphisms will be *-isomorphisms—that is an isomorphism $\alpha$ such that $\alpha(A^{\dagger})=\alpha(A)^{\dagger}$ for all operators $A$. All such automorphisms of this kind between matrix algebras can be expressed as a conjugation by a unitary.) To find a QCA $V$ which implements this isomorphism simultaneously for all sites, we need to match the algebras $\mathcal{O}_{k}$ to $\mathcal{T}_{j(k)}$ locally and bijectively. This can be phrased as a graph problem, and Hall’s marriage theorem [122, 123] implies that this problem always has a solution.
@@ A.3 Non-uniqueness
where $\pi$ is a permutation of the simultaneous eigenstates of the $\{\sigma^{z}_{k}\}_{k\in K}$ operators and ${\vec{m}}_{\alpha}\in{\mathbb{Z}}^{n}$ are vectors of integers. (Such $P({\vec{\theta}})$ are generalized permutation matrices in a frequency lattice picture [31, 99].) Any $P$ of this form defines an identical unitary evolution (35) with
@@ Appendix B The $\Omega$-spectrum property
For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial. In positive degree, we propose that the swindle map $S$ (Eq. (8)) provides the homotopy equivalence. Thus, we propose there is a generalized cohomology theory
@@ B.2 Deformation to the swindle
Figure 4: The loop of finite depth circuits $\gamma(t)\otimes S(v_{0}^{\gamma},s)^{\dagger}$ is deformed to the constant loop at the identity as a function of $s\in[0,1]$. (a) At $s=0$, the circuit consists of two layers. Between $s=0$ and $s=1/2$ the gates $u^{\dagger}_{x,x+1}(v_{0}^{\gamma},t)$ in the top layer (red) are swapped to the bottom layer while having their order of multiplication reversed. (b) The gate $u^{\dagger}_{0,1}(v_{0}^{\gamma},t)w_{0,1}(t)$ ($u^{\dagger}_{-1,0}(v_{0}^{\gamma},t)w_{-1,0}(t)$) acts as the identity on site 00 for $t\in\{0,1\}$. It can thus be swapped (over $s\in[\tfrac{1}{2},\tfrac{3}{4}]$) so that it acts on site $1$ ($-1$) and its ancilla while ensuring that the whole circuit is still a loop. (c) Between $s=\tfrac{3}{4}$ and $s=1$ the circuit (red) is “rolled up”, such that it acts as the identity (gray dots) in any finite region.
@@
This rolling up procedure can be iterated. At $s=\frac{3}{4}$, we see that the gate $w^{\prime}_{1,2}(t)w^{\prime}_{1^{\prime},1}(t)$ is the identity on sites $1$ and $1^{\prime}$ when $t=1$. Thus, by adding ancillae to site $2$, this gate can also be rolled up within $s\in[\tfrac{3}{4},\tfrac{7}{8}]$ (Fig. 4(c)). We continue rolling up sites $x$ and $-x$ within $s\in[1-2^{-x-1},1-2^{-x-2}]$, such that, as $s\to 1$, we have $E(\gamma,t,s)\to\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ on any finite patch of the lattice. Technically, this makes $E(\gamma,t,s)$ continuous in the strong operator topology.
The deformation is completed by adding a parallel layer of swap gates which takes $\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ to $S(v_{0}^{\gamma},t)\otimes\mathbbm{1}\in\mathrm{im}(S)$. If the circuit decomposition Eq. (61) gives $w_{x,x+1}=u_{x,x+1}(v)$ (as it does when using the $\mathrm{Tr}_{(x,\infty)\times{\mathbb{Z}}^{d}}{H(t)}$ construction of Appendix B.1) then $E$ is a deformation retract, and Eq. (51) holds.
@@ C.1 Connected components
For one-dimensional QCA, the complete classification (that is, $\pi_{0}(\mathrm{QCA}_{1})$) was obtained by Ref. [58], which also explicitly defined an invariant which diagnoses which connected component a given QCA belongs to. All one-dimensional QCA are path connected to a shift—a QCA which acts as a translation on some subset of the degrees of freedom. For example, in a chain of qubits, each QCA is connected to a QCA which translates all qubits by $r$ sites either right or left. For qubit chains, QCA are classified by the group ${\mathbb{Z}}$.
@@ C.2.2 Stable invariants
These are no longer integer invariants, but rather rational invariants (modulo 1 for $n=1$). Then Eq. (78) shows that $\tilde{W}_{n}[V\otimes Q]=\tilde{W}_{n}[V]+\tilde{W}_{n}[Q]$.

### v2

@@ Abstract
This Letter systematizes the study of ALT phases in strongly interacting systems. Specifically, we focus on localized topological phases satisfying a condition we refer to as short-ranged entanglement. We will argue that instead of eigenstate order, the right way to characterize all topological localized phases with short-ranged entanglement is via quantum cellular automata (QCA) [58, 59, 60, 56]—the locality preserving unitaries which map physical degrees of freedom to local integrals of motion (LIOMs) [Fig. 1(a)].
As QCA in spin systems have been classified in one [58] and two [61] dimensions, this correspondence allows for a classification of ALT phases in low dimensional systems. The same formalism extends to periodically (Floquet [62]) and quasiperiodically driven ALT phases [31], which are classified by loops or tori of QCA, respectively. Thus, we can treat static and $n$-tone-driven ALT phases in $d$ dimensions on the same footing. We also considerably develop the study of the topology of QCA, which allows us to classify all ALT phases with $d-n\leq 2$ [Fig. 1(b)].
@@
We say two MBL Hamiltonians $H_{0}$ and $H_{1}$ are in the same localized topological phase if there is a continuous path of MBL Hamiltonians $H_{t}$ ($t\in[0,1]$) which connects them. Ignoring long-range resonances in the spectrum of $H_{t}$, it is possible to construct a continuous path of locality preserving unitaries $V_{t}$ (that is, $V_{t}AV_{t}^{\dagger}$ is localized near the support of $A$ for all operators $A$) such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$ [84, 85, 86, 87, 88, 89]. Indeed, this unitary simply follows the adiabatic evolution of the many-body eigenstates. Including long-range resonances prevents an adiabatic construction of $V_{t}$ [90], but it is expected that sufficiently dilute resonances can be incorporated such that there still exists a $V_{1}$ that conjugates LIOMs of $H_{0}$ into those of $H_{1}$ and is continuously connected to the identity [9]. Thus, if $H_{0}$ and $H_{1}$ are in the same localized phase, their LIOMs are related by a continuous path of locality-preserving unitary operators. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@
$\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}\implies\pi_{n-1}(\mathrm{QCA}_{d-1})\cong\pi_{n}(\mathrm{QCA}_{d}).$ (5)
The base point for the loops in $\Omega\mathrm{QCA}_{d}$ is the identity.
@@
Equation (5) is interpreted through a picture of pumping [95, 104, 103]. Taking a loop $\gamma\in\Omega\mathrm{QCA}_{d}$ and restricting it to open boundary conditions, the action in the bulk must still be trivial at the end of the loop, due to the finite range of the QCA. However, on the $(d-1)$-dimensional edge, there may be a nontrivial action by some $v\in\mathrm{QCA}_{d-1}$ [23]. We say that $\gamma$ pumps $v$ across the system, and Eq. (5) hypothesizes that the homotopy structure of the loop space $\Omega\mathrm{QCA}_{d}$ is completely characterized by the pumped QCA $v$.
@@
The conjectured classification of symmetry-enriched invertible states can be stated in terms of just the space $\mathrm{inv}_{d}$ and the abstract symmetry group $G$ [96, 66, 95, 102, 113, 103, 104]. However, the classification of symmetric QCA depends not just on $G$, but on the specific representation $\rho$ of the symmetry action [93, 114]. We compute the classification of symmetric QCA, and thus SALT phases, with $d-n\leq 1$ and any on-site unitary representation of any group in Appendix D. The classification of symmetric QCA closely mirrors Tab. 1, except that it includes additional subgroups in $d-n\geq 0$ [given by $H^{d-n+1}(G;U(1))$ for finite $G$] related to nontrivial SPT states or pumps thereof; moreover, rational numbers $p/q\in{\mathbb{Q}}$ in Tab. 1 are replaced with formal ratios of representations, $\rho/\sigma$. Addition and multiplication for $\rho/\sigma$ is defined through the same rules as for ratios of integers, but replacing “$+$” with $\oplus$” and “$\times$” with $\otimes$”. In an appropriate coincidence, this mathematical structure is known as an algebraic localization.
To define the algebraic localization precisely, we must specify the set of ancillae representations,
@@
which is closed under tensor product—if $\rho_{1}$ and $\rho_{2}$ can be added then $\rho_{1}\otimes\rho_{2}$ corresponds to adding them both. The algebraic localization with denominators in $\mathcal{R}$ is the quotient
@@
where $R_{G}$ is the representation ring of $G$, and $\simeq$” is unitary equivalence of representations. Equivalence classes of pairs $(\rho,\sigma)$ are denoted $\rho/\sigma$.
@@
Models.—While the swindle map [Eq. (8), Fig. 2] was introduced as an abstract tool, it also defines a class of soluble models of $n$-tone-driven (S)ALT phases given any $(n-1)$-parameter family of QCA in one dimension lower. As these models are finite depth circuits, they are also well suited for realization in quantum simulator experiments [63, 64, 65] (Fig. 3).
@@
The general relationship between (S)ALT phases and QCA we have demonstrated, as well as the $\Omega$-spectrum property which relates static (S)ALT phases to driven phases, extends to fermionic systems essentially without modification (Appendix A-B[56]. The $\Omega$-spectrum structure of the classification can also be exploited to classify SALT phases with lattice symmetries [116, 103, 104]. The more challenging direction of generalization is to include topological order (long-ranged entanglement), where the correspondence with QCA breaks down [20]. This case is of particular interest given recent proposals of using driven topologically ordered localized models to perform quantum computation [117, 118, 119, 120]. The notion of measurement QCA introduced by Ref. [121] may provide a route towards classifying localized phases with topological order, which we will explore in future work.
The authors are particularly grateful to R. Kobayashi for collaboration during the early stages of this project. The authors also thank M. Barkeshli, A. Chandran, L. Fidkowski, S. Pace, R. Thorngren, Y.-T. Tu, B. Ware, and C. Zhang for useful discussions. This work is supported by the Laboratory for Physical Sciences (DML). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities (DVE).
@@ A.1 Setup
[where $B_{r}(j)$ is the ball of radius $r$ around $j$] such that $[\tau^{z}_{j},H]=[\tau^{z}_{j},\tau^{z}_{j^{\prime}}]=0$. Further, the set of LIOMs must be maximal, such that a local operator $A$ commutes with all LIOMs only if $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$.
@@
This definition should be compared to [56, Definition II.3]. One can also compare this notion of “short-ranged entanglement” for Hamiltonians to the usual notion of short-ranged entanglement for states. A state $|\psi\rangle$ may be said to be short-range entangled if, for any other state $|\psi^{\prime}\rangle$ which differs from $|\psi\rangle$ only in some ball $B_{r}(x)$ [$\langle\psi|A|\psi\rangle=\langle\psi^{\prime}|A|\psi^{\prime}\rangle$ for any local $A$ not supported in $B_{r}(x)$] there is a local unitary $\tau^{x}_{\psi\psi^{\prime}}$ supported in $B_{r}(x)$ such that $\tau^{x}_{\psi\psi^{\prime}}|\psi^{\prime}\rangle=|\psi\rangle$. Any eigenstate of an $H$ satisfying Definition A.2 is short-range entangled in this sense (the required $\tau^{x}_{\psi\psi^{\prime}}$ can be constructed from the $\tau^{x}_{j}$ operators), but the converse—that $\tau^{x}_{j}$ operators exist for any MBL Hamiltonian $H$ whose eigenstates are all short-range entangled—is not obvious.
@@ A.2 Existence
is isomorphic to a matrix algebra. This implies that there is an isomorphism between $\mathcal{T}_{j}$ and $\mathcal{O}_{k}$ whenever they have the same dimension. [All isomorphisms will be *-isomorphisms—that is an isomorphism $\alpha$ such that $\alpha(A^{\dagger})=\alpha(A)^{\dagger}$ for all operators $A$. All such automorphisms of this kind between matrix algebras can be expressed as a conjugation by a unitary.] To find a QCA $V$ which implements this isomorphism simultaneously for all sites, we need to match the algebras $\mathcal{O}_{k}$ to $\mathcal{T}_{j(k)}$ locally and bijectively. This can be phrased as a graph problem, and Hall’s marriage theorem [122, 123] implies that this problem always has a solution.
@@ A.3 Non-uniqueness
where $\pi$ is a permutation of the simultaneous eigenstates of the $\{\sigma^{z}_{k}\}_{k\in K}$ operators and ${\vec{m}}_{\alpha}\in{\mathbb{Z}}^{n}$ are vectors of integers. [Such $P({\vec{\theta}})$ are generalized permutation matrices in a frequency lattice picture [31, 99].] Any $P$ of this form defines an identical unitary evolution (35) with
@@ Appendix B The $\Omega$-spectrum property
For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial. In positive degree, we propose that the swindle map $S$ [Eq. (8)] provides the homotopy equivalence. Thus, we propose there is a generalized cohomology theory
@@ B.2 Deformation to the swindle
Figure 4: The loop of finite depth circuits $\gamma(t)\otimes S(v_{0}^{\gamma},s)^{\dagger}$ is deformed to the constant loop at the identity as a function of $s\in[0,1]$. (a) At $s=0$, the circuit consists of two layers. Between $s=0$ and $s=1/2$ the gates $u^{\dagger}_{x,x+1}(v_{0}^{\gamma},t)$ in the top layer (red) are swapped to the bottom layer while having their order of multiplication reversed. (b) The gate $u^{\dagger}_{0,1}(v_{0}^{\gamma},t)w_{0,1}(t)$ [$u^{\dagger}_{-1,0}(v_{0}^{\gamma},t)w_{-1,0}(t)$] acts as the identity on site 00 for $t\in\{0,1\}$. It can thus be swapped (over $s\in[\tfrac{1}{2},\tfrac{3}{4}]$) so that it acts on site $1$ [$-1$] and its ancilla while ensuring that the whole circuit is still a loop. (c) Between $s=\tfrac{3}{4}$ and $s=1$ the circuit (red) is “rolled up”, such that it acts as the identity (gray dots) in any finite region.
@@
This rolling up procedure can be iterated. At $s=\frac{3}{4}$, we see that the gate $w^{\prime}_{1,2}(t)w^{\prime}_{1^{\prime},1}(t)$ is the identity on sites $1$ and $1^{\prime}$ when $t=1$. Thus, by adding ancillae to site $2$, this gate can also be rolled up within $s\in[\tfrac{3}{4},\tfrac{7}{8}]$ [Fig. 4(c)]. We continue rolling up sites $x$ and $-x$ within $s\in[1-2^{-x-1},1-2^{-x-2}]$, such that, as $s\to 1$, we have $E(\gamma,t,s)\to\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ on any finite patch of the lattice. Technically, this makes $E(\gamma,t,s)$ continuous in the strong operator topology.
The deformation is completed by adding a parallel layer of swap gates which takes $\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ to $S(v_{0}^{\gamma},t)\otimes\mathbbm{1}\in\mathrm{im}(S)$. If the circuit decomposition Eq. (61) gives $w_{x,x+1}=u_{x,x+1}(v)$ [as it does when using the $\mathrm{Tr}_{(x,\infty)\times{\mathbb{Z}}^{d}}{H(t)}$ construction of Appendix B.1] then $E$ is a deformation retract, and Eq. (51) holds.
@@ C.1 Connected components
For one-dimensional QCA, the complete classification [that is, $\pi_{0}(\mathrm{QCA}_{1})$] was obtained by Ref. [58], which also explicitly defined an invariant which diagnoses which connected component a given QCA belongs to. All one-dimensional QCA are path connected to a shift—a QCA which acts as a translation on some subset of the degrees of freedom. For example, in a chain of qubits, each QCA is connected to a QCA which translates all qubits by $r$ sites either right or left. For qubit chains, QCA are classified by the group ${\mathbb{Z}}$.
@@ C.2.2 Stable invariants
These are no longer integer invariants, but rather rational invariants [modulo 1 for $n=1$]. Then Eq. (78) shows that $\tilde{W}_{n}[V\otimes Q]=\tilde{W}_{n}[V]+\tilde{W}_{n}[Q]$.